### v1

Abstract
Bilayers (soft substrates coated with stiff films) are commonly found in nature with examples including skin tissue, vesicles, or organ membranes. They exhibit various types of instabilities when subjected to compression, depending on the contrast in material properties between the two components. We present wrinkling instabilities for 3D hyperelastic bilayer systems, including auxetics (materials with negative Poisson’s ratio), under uni-axial tension. In tension, a soft bilayer can experience large lateral contraction, and we find that with an adequate contrast in the Poisson ratios, compressive stresses may develop and generate wrinkles aligned with the tensile direction. We rely on an analytic modelling of the phenomenon, and validate it with a user-defined Python script with periodic boundary conditions and constitutive relation implementation in advanced Finite Element simulations. Our findings reveal that wrinkles are observed when the Poisson ratio of the substrate is greater than that of the film. As the two Poisson ratios converge to a common value, the critical stretch of instability shoots up rapidly, and the wrinkling disappears. We also confirm these results by asymptotic analysis. This wrinkling analysis has significant potential in controlling surface patterns of auxetic skin grafts and hydrogel organ patches under mechanical loads. Moreover, the asymptotic expressions in this work can be used under finite strain for buckling-based metrology applications.
keywords:
[…]
1 Introduction
Auxetics, materials with negative Poisson’s ratio, expand in all directions under uni-axial tension. For 3D isotropic materials with auxetic behavior, the theoretical value of Poisson’s ratio ranges between -1 and 0.5 [1]. Thanks to rapid advancement in additive and subtractive manufacturing techniques [2, 3], along with extensive research on negative Poisson ratio materials, auxetics have shown promising applications in various fields [4, 5, 6, 7, 8, 9].
Compliant substrates coated with thin-layered stiff films (bilayers) are commonly found in nature; for example, skin tissue consists of a thin, stiff epidermal layer attached to a thick, soft dermis. When subject to mechanical loads, bilayer systems can exhibit surface patterns through wrinkles. This instability phenomenon has found a broad range of applications, in optical sensors [10], novel flexible electronics [11, 12, 13], tunable phase gratings [14], buckling-based metrology [15, 16], surface wetting [17], and for buckling-related applications in soft matter [18].
The formation of these surface patterns is controlled by different parameters, such as the contrast in material properties [19, 20], differential growth [21], film-to-substrate thickness ratio [22, 23], initial imperfections [24], substrate nonlinearity [25, 26], interfacial mechanics [27, 28], etc.
A survey of the literature shows that wrinkles appearing in tension are seldom studied beyond the linear-elastic framework [29], although they have been established experimentally under large strains [30, 15], see Fig. 1. In this work, we perform a linear buckling analysis of 3D hyperelastic thin stiff films on semi-infinite compliant substrates (including auxetics) in tension and large strains, in contrast to the numerous existing studies concerned with compression.
We first use the buckling analysis available in ABAQUS [31], which works up to a point, beyond which we have to rely on a semi-analytical approach with Mathematica [32]. We find that wrinkles may be generated in bilayers subjected to uni-axial tension. Main and original findings include (i) Deriving asymptotic expressions for critical stretch ratios and critical wavenumbers, valid in finite strains and useful for buckling-based metrology applications; (ii) Showing that the Poisson ratio of the substrate must be greater than that of the film for wrinkles to appear; (iii) Establishing that wrinkles disappear when the Poisson ratios of film and substrate converge to a common value, and that (iv) The wrinkling wavelength is high (low) when the Poisson ratios of film and substrate are close (far).
2 Results
2.1 3D bilayer system in tension
Figure 1: (a) Undeformed and (b) deformed configurations of a 3D bilayer system with periodic boundary conditions in the lateral directions. The system is infinite in the X- and Z-directions and is under uni-axial tension along the X-direction. The analysis shows that eventually, (b) wrinkles develop parallel to the X-direction, provided there is enough contrast between the constitutive parameters of the materials. (c) Image obtained using Scanning Electron Microscope (SEM) on a natural rubber substrate coated with a gold film, when elongated to 50% strain at room temperature [30]. An inset with an arrow shows the direction of elongation. Bright and dark bands represent the buckled gold film layer and the rubber showing through cracks in the stretched film, respectively. (d) Optical micrograph of a Polystyrene (PS) film layer on a silicon wafer with decreasing thickness from top to bottom is shown on the left panel. The corresponding wrinkles exhibit decreasing wavelengths on the right panel, when the PS films are attached to a Polydimethylsiloxane (PDMS) substrate and elongated to induced buckling [15].
We consider 3D bilayer systems with a thin stiff film perfectly bonded to a semi-infinite compliant substrate, see Fig. 1. Under a uniaxial tension applied along the X-direction, the bilayer is expected to eventually develop wrinkles aligned with that direction, as shown in Fig. 1(b), provided compressive stresses develop along the Y-direction because of a sufficient Poisson ratio contrast.
Figure 2: Comparison of wrinkling profiles in thin conventional and auxetic sheets. Top row: Thin acetate sheets under uni-axial tension, with (a) non-auxetic micro-structural patterns and (b) auxetic micro-structural patterns [33]. For conventional sheets, wrinkles develop at the center of the sheet. For auxetic sheets, they appear near the clamped edges. Bottom row: Buckling profiles (according to the magnitude of the displacement field) obtained with ABAQUS and the Blatz-Ko model Eq. 1. Here $\alpha=0.4$, $\mu=0.53$ GPa, and $\nu=0.38,-0.2$ in (c) and (d), respectively.
Here we use the Blatz-Ko strain energy function [34] for both materials:
$W=c_{1}\left(I_{1}-3+\frac{1}{\beta}\left(I_{3}^{-\beta}-1\right)\right)\\ +c_{2}\left(\frac{I_{2}}{I_{3}}-3+\frac{1}{\beta}\left(I_{3}^{\beta}-1\right)\right),$ (1)
where $c_{1}\!=\!\alpha\frac{\mu}{2}$, $c_{2}\!=\!(1-\alpha)\frac{\mu}{2}$, $\beta=\frac{\nu}{1-2\nu}$, $\mathbf{F}$ is the deformation gradient, and $I_{1}=\text{tr}(\mathbf{FF}^{T})$, $I_{2}=I_{3}\text{tr}[(\mathbf{FF}^{T})^{-1}$], $I_{3}=\det(\mathbf{FF}^{T})$ are three strain invariants. Also, the material constants are the non-dimensional parameter $10<\alpha<$ 1, the initial shear modulus $\mu>0$, and the Poisson ratio $-1<\nu\leq 1/2$. We use the letters f and s for film and substrate, respectively. For the simulations, $\alpha_{f}=\alpha_{s}=0.4$ and $\mu_{f}/\mu_{s}=30$ (similar observations to those detailed in this study can be made for other values).
Auxetic materials are highly compressible due to their negative Poisson ratio (thus far away from the incompressibility limit of a Poisson ratio equal to 1/2). Blatz-Ko strain energy functions with a negative Poisson ratio can thus be used to model auxetics, although with certain limitations [35]. For example, Ciambella and Saccomandi [36] use the Blatz-Ko strain energy to capture the experiments of Choi and Lakes [37] on auxetic materials. The model is also adequate to capture the experimental behaviors of membranes (no substrate) under tension, from auxetic to conventional Poisson ratios, see Fig. 2.
Now, using the Blatz-Ko model for 3D bilayer systems under large uni-axial elongation, we find that wrinkles are observed only when the Poisson ratio of the film ($\nu_{f}$) is smaller than that of the substrate ($\nu_{s}$), as summarised in Fig. 3. This result recovers the work by Nikravesh et al. [38] conducted for linear-elastic bilayer systems.
Figure 3: Wrinkling condition according to the contrast between the Poisson ratios of film and substrate.
2.2 Incompressible substrate
Figure 4: Numerical simulations using Blatz-Ko material model (Eq. 1). Variations of the critical stretch of wrinkling $\lambda_{c}$ and corresponding critical wavenumber measure $k_{c}H$ with the Poisson ratio of one layer ($k_{c}$ is the critical wavenumber and $H$ is the film initial thickness). (a-b): The substrate is quasi-incompressible ($\nu_{s}=0.495$). (c-d): The film is highly auxetic ($\nu_{f}=-0.95$). Results from ABAQUS: black dots, results from Mathematica: solid line with square markers.
First, we assume that the substrate is quasi-incompressible ($\nu_{s}=0.495$) while for the film, $\nu_{f}\in(-0.95,0.495)$.
Here the Poisson ratio of the substrate is greater than that of film, and under uniaxial tension, the film expands more than the substrate along the Z-direction, resulting in compressive stresses and the formation of wrinkles parallel to the X-direction.
In Fig. 4(a-b), the values of the critical stretch of wrinkling $\lambda_{c}$ and of the corresponding critical wavenumber $k_{c}$ are plotted against $\nu_{f}$. From Fig. 4a, we observe that the critical stretch ratio values predicted from the semi-analytical analysis (Mathematica) and the FE buckling analysis ABAQUS match well, as long as $\nu_{f}\leq 0.35$. Beyond this value, ABAQUS ceases to predict wrinkles in the desired direction and gives negative eigenvalues, suggesting the load direction has to be reversed to obtain wrinkles, which is unphysical.
An auxetic film ($\nu_{f}<0$) expands in all directions under tension while the incompressible substrate contracts along the Z direction, leading to compressive stress in the film. Hence only low values of the critical stretch are required for the wrinkles to occur. As the Poisson ratio of the film increases and moves closer to that of the substrate, the values of the critical stretch shoot up sharply, supporting our findings in Fig. 3.
2.3 Highly auxetic film
Now we take the film to be highly auxetic, with Poisson ratio $\nu_{f}=-0.95$, while for the substrate, $\nu_{s}\in(-0.95,0.495]$.
Fig. 4(c-d) show the variations of $\lambda_{c}$ and $k_{c}$ with $\nu_{s}$. Again, we find good agreement between analysis and simulations for a certain range, when $-0.7\leq\nu_{s}\leq 0.495$. For $\nu_{\textbf{s}}<-0.8$, ABAQUS stops providing meaningful predictions. In general, the buckling occurs early, except as $\nu_{s}\to\nu_{f}$ when it increases dramatically because in that limit, both film and substrate experience the same transverse contraction so that no compressive stress nor wrinkles develop. Similarly to Fig. 4b, Fig. 4d shows that the higher the Poisson ratio contrast is, the larger the wavelength of the wrinkles is.
In summary, we found that to ensure early wrinkling in tension, a large contrast in Poisson’s ratio is required, for example by taking one material to be auxetic and the other non-auxetic.
2.4 Asymptotic solution: Compressible neo–Hookean bilayer system
The following critical strain and wavenumber expressions:
$\varepsilon_{c}=\frac{1}{4(\nu_{s}-\nu_{f})}\left(3\frac{\mu_{s}}{\mu_{f}}\frac{1-\nu_{f}}{1-\nu_{s}}\right)^{2/3},$
$k_{c}h=\left(3\frac{\mu_{s}}{\mu_{f}}\frac{1-\nu_{f}}{1-\nu_{s}}\right)^{1/3},$ (2)
have been derived by [30, 39, 38] for linearly elastic bilayers under uniaxial tension. These expressions are valid under the plane-strain approximation. Here, $h$ is the current thickness of the film, $\varepsilon_{c}$ is the critical strain, and $k_{c}$ is the critical wavenumber.
Figure 5: Numerical simulations for a bilayer made of compressible neo-Hookean materials ($\alpha_{s}=\alpha_{f}=1$ in Eq. 1). The substrate is quasi-incompressible ($\nu_{s}=0.495$). (a-b): Variations of the critical stretch of wrinkling $\lambda_{c}$ and corresponding critical wavenumber measure $k_{c}h$ ($k_{c}$: critical wavenumber, $h$: current film thickness) with the Poisson ratio of film layer $\nu_{f}$. ABAQUS: black dots, Mathematica: solid line with square markers, Asymptotic expressions (Eqs. 3 to 4): star markers, Linear-elastic expressions (Eq. 2): dash-dotted line.
However, we note that the plane-strain assumption is not valid for a bilayer system under uni-axial tension. Here, we follow Cai and Fu [40, 41] and assume that the shear modulus ratio $r={\mu_{s}}/{\mu_{f}}$ is of order $(k_{c}h)^{3}$ to derive asymptotic expansions for the critical strain and wavenumber (for simplicity, we take $\alpha_{s}=\alpha_{f}=1$ in Eq. 1). We find
$\lambda_{c}=1$ $+\frac{\left[(1-\nu_{f})(1-\nu_{s})\right]^{2/3}}{(\nu_{s}-\nu_{f})\left(8\nu_{s}/3-2\right)^{2/3}}r^{2/3}$
$+\frac{(1-\nu_{f})(2\nu_{s}-1)}{(\nu_{s}-\nu_{f})(4\nu_{s}-3)}r$
$+c_{4}r^{4/3}+c_{5}r^{5/3}+c_{6}r^{2}+\pazocal{O}(r^{7/3}),$ (3)
and
$k_{c}h=$ $\left[\frac{4(1-\nu_{f})(1-\nu_{s})}{1-4\nu_{s}/3}\right]^{1/3}r^{1/3}$
$+d_{3}r+d_{4}r^{4/3}+d_{5}r^{5/3}+\pazocal{O}(r^{2}),$ (4)
where the coefficients $c_{4},\cdots,c_{6},d_{3},\cdots,d_{5}$ are shown in Appendix B: Coefficients in asymptotic expressions.
[…]
In Fig. 6(a-b), we plot the variations of $\lambda_{c}$ and $k_{c}h$ with the Poisson ratio of the film ($0.450<\nu_{f}<$ 0.45) for different values of shear moduli contrast $\mu_{f}/\mu_{s}$ using Eq. 3-Eq. 4, because they give curves instantly compared to long computations.
Figure 6: Using the asymptotic expressions Eq. 3-Eq. 4 for compressible neo-Hookean materials ($\alpha_{s}=\alpha_{f}=1$ in Eq. 1). The substrate is quasi-incompressible ($\nu_{s}=0.495$). (a-b): Variations of the critical stretch of wrinkling $\lambda_{c}$ and corresponding critical wavenumber measure $k_{c}h$ with the Poisson ratio of film layer ($k_{c}$ is the critical wavenumber and $h$ is the deformed film thickness) and for contrast in shear moduli between the layers $(\nicefrac{{\mu_{f}}}{{\mu_{s}}}=\{50,100,500,1000\})$.
3 Methods
3.1 Semi-analytical treatment
We use Eq. 1 to determine the principal components of the Cauchy stress $\mathbf{T}=J^{-1}\mathbf{F}\left(\partial W/\partial\mathbf{F}\right)$ as
$T_{i}=E\frac{J^{-(2\alpha+1)}\left(-\alpha\lambda_{i}^{2}+(\alpha-1+\alpha\lambda_{i}^{4})J^{2\alpha}-(\alpha-1)\lambda_{i}^{2}J^{4\alpha}\right)}{2\lambda_{i}^{2}\left(1+\nu\right)},$ (5)
where the $\lambda_{i}$ are the principal stretch ratios and $J=\det\mathbf{F}$.
The substrate is under uniaxial tension along X, so that $T^{s}_{22}=T^{s}_{33}=0$, which gives the following deformation gradient,
$\mathbf{F}^{{s}}=\text{diag}\left(\lambda_{1},\lambda_{1}^{-\nu_{{s}}},\lambda_{1}^{-\nu_{{s}}}\right).$ (6)
The film and substrate are perfectly bonded, so that $\lambda_{1}^{f}=\lambda_{1}^{s}$, $\lambda_{3}^{f}=\lambda_{3}^{s}$, and $T^{f}_{22}=T^{s}_{22}=0$. We then find that
$\mathbf{F}^{{f}}=\text{diag}\left(\lambda_{1},\lambda_{1}^{\frac{\nu_{{f}}(\nu_{{s}}-1)}{(1-\nu_{{f}})}},\lambda_{1}^{-\nu_{{s}}}\right).$ (7)
To find the critical state of buckling, a small-amplitude mechanical displacement $\mathbf{u}$ is superimposed on the finite deformations. The incremental equations of equilibrium read
$\pazocal{A}_{0jilm}^{{s}}u_{m,lj}^{{s}}=0,$ $-\infty<y<0,$
$\pazocal{A}_{0jilm}^{{f}}u_{m,lj}^{{f}}=0,$ $0<y<h,$ (8)
where the commas denote differentiation with respect to the coordinates, $h$ is the current thickness of the film, and $\pazocal{A}_{0}$ is the fourth-order tensor of the instantaneous elastic moduli, with components
$\pazocal{A}_{0jilm}=J^{-1}F_{j\alpha}\frac{\partial^{2}W}{\partial F_{i\alpha}\partial F_{m\beta}}F_{l\beta}.$ (9)
We look for solutions of the forms:
$u_{z}=e^{sy}\sin(kz),\qquad u_{y}=e^{sy}\cos(kz),$ (10)
where $s$ is the attenuation coefficient and $k$ is the wavenumber of the sinusoidal wrinkles. Substitution into Eq. 8 leads to an eigenproblem, with characteristic equation a bicubic in $s$. In the film, the general solution is of the form
$\begin{split}u_{z}^{{f}}&=\Big{(}\sum_{i=1}^{4}\pazocal{V}_{i}e^{s_{i}y}\Big{)}\sin(kz),\quad u_{y}^{{f}}=\Big{(}\sum_{i=1}^{4}\pazocal{V}_{i}e^{s_{i}y}\Big{)}\cos(kz),\end{split}$ (11)
where $s_{1},\cdots,s_{4}$ are the eigenvalues, and $\pazocal{V}_{1},\cdots,\pazocal{V}_{4}$ are constants. In the substrate, the stretch ratios along the Y- and Z- directions are equal, and $s=1$ is a repeated eigenvalue. The other repeated root, $s=-1$ is discarded to enforce decay. There, the solution is thus of the form
$\begin{split}u_{z}^{{s}}&=(\pazocal{U}_{1}+\pazocal{U}_{2}y)e^{y}\sin(kz),\qquad u_{y}^{{s}}=(\pazocal{U}_{1}+\pazocal{U}_{2}y)e^{y}\cos(kz),\end{split}$ (12)
where $\pazocal{U}_{1},\pazocal{U}_{2}$ are constants.
By applying the traction-free boundary conditions
$\pazocal{A}_{02ilm}^{{f}}u_{m,l}^{{f}}=0,\qquad y=h,$ (13a)
and the continuity conditions
$\pazocal{A}_{02ilm}^{{f}}u_{m,l}^{{f}}=\pazocal{A}_{02ilm}^{{s}}u_{m,l}^{{s}},\qquad u_{i}^{{f}}=u_{i}^{{s}},\qquad y=0,$ (13b)
we obtain six homogeneous equations for $\{{\pazocal{U}_{1},\pazocal{U}_{2},\pazocal{V}_{1},\cdots,\pazocal{V}_{4}\}}$. The bifurcation condition is then given by equating the determinant of a $6\times 6$ coefficient matrix to 0.
Figure 7: (a) Undeformed and (b) deformed configurations of auxetic stiff film with orthogonal-oval shaped patterns bonded to a compliant substrate. The base materials for film and substrate are incompressible. Periodic boundary conditions are applied on the extreme faces of the domain along X and Z and the system is subjected to uni-axial tension along X. In (b), we observe that the wrinkles are generated parallel to the direction of tension. The critical strain for buckling is $\varepsilon_{c}\simeq 0.274.$
3.2 Finite Element simulations: 3D models
We first checked the accuracy of our periodic boundary conditions (PBCs) code (Python script) to reproduce known results for wrinkling under plane-strain compression. Hence, we performed a linear buckling analysis in ABAQUS on an incompressible neo-Hookean bilayer periodic along the X-direction, see Appendix A: Validation of PBCs code – 2D incompressible neo-Hookean bilayer model, and recovered the results of [42].
We then perform a series of finite-element simulations for 3D Blatz-Ko model bilayers under uni-axial tension. The dimensions (X, Y, Z) of the bilayer structure are $15\times 45\times 5.684$ units with ${h_{{s}}}/{h_{{f}}}=299$ when the substrate is incompressible, and $15\times 80\times 4.504$ units with ${h_{{s}}}/{h_{{f}}}=399$ when the film is highly auxetic.
The dimensions along the Z-direction are integer multiples of the wavelength (calculated using semi-analytical results from Mathematica) when $(\nu_{f},\nu_{s})=\{0.3$, 0.495}0.495\}0.495 } and $\{-0.95,-0.8\}$. For the other $(\nu_{f},\nu_{s})$ pairs, we noticed that small variations in the depth of the structure did not notably alter the values of critical stretch (changes in the third decimal point). To save computational time (without compromising the efficiency of the solutions), we kept the dimensions $15\times 45\times 5.684$ units and $15\times 80\times 4.504$ units when the substrate is incompressible (Section 2.2) and when the film is highly auxetic (Section 2.3).
For the neo-Hookean bilayers ($\alpha_{f}=\alpha_{s}=1$), the dimensions are $6\times 45\times 1.85$ units with ${h_{{s}}}/{h_{{f}}}=299$.
We apply PBCs on the extreme faces in the X- and Z-directions of the domain using a user-defined Python script file. Both the film and substrate are modelled using a UHYPER subroutine with the Blatz-Ko strain energy function. We use a 20-node brick (hexahedral) element with quadratic interpolation and reduced integration (C3D20R) for both film and substrate; for the limiting case of incompressibility, we used a C3D20RH element. For the Blatz-Ko bilayers, 34,200 and 17,130 mesh elements are used for the cases of an incompressible substrate and a highly auxetic film, respectively. For the neo-Hookean bilayer system, 3,240 mesh elements are used.
We find that the numerical results obtained using different approaches in ABAQUS and Mathematica software match well for a wide range of material parameters, see Fig. 4. When $\alpha_{f}=\alpha_{s}=1$ (compressible neo-Hookean bilayers), the results using asymptotic expressions agree well with the numerical simulations, see Fig. 5.
4 Conclusions
We investigated the possibility of harnessing wrinkles parallel to the direction of applied tension in 3D isotropic compressible bilayers subject to large elongations. We paid particular attention to the cases when the substrate is quasi-incompressible and when the film is highly auxetic.
We used a semi-analytical approach with Mathematica to predict the onset of wrinkling, and user-defined Python scripts to apply periodic boundary conditions and UHYPER subroutine for Blatz-Ko material models to perform linear buckling analysis in Finite Element simulations with ABAQUS.
For compressible neo–Hookean bilayer systems, we derived asymptotic expressions for critical stretch ratio and critical wavenumber, which can be used under finite strains to determine the Young modulus of the film layer for buckling-based metrology applications.
We found that wrinkles can be obtained only when the Poisson ratio of the substrate is greater than that of the film. When the Poisson ratios of film and substrate converge to a common value, the critical stretch ratio shoots up sharply and the wavelength of wrinkles is high. In the limit, the wrinkles are not present because there are no compressive stresses developing when the lateral expansions are the same for the film and substrate. Through multiple simulations we showed that by varying the material properties, we can harness or delay the onset of wrinkles.
Some of the limitations of our work include the consideration of isotropic strain energy function for auxetic materials, and deformation-independent material properties. Functional-grading of auxetics could also be explored with the methods presented in this study, see some preliminary works on harnessing instabilities in functionally-graded auxetic materials using tension-field theory [43, 44] and their applications [45, 46, 47].
Another area of interest concerns instabilities in materials with auxetic patterns. Indeed, auxetic properties can be obtained at a continuum level with a careful design of holes or voids at the micro-scale [6, 48]. For example, in Fig. 7 we performed linear buckling analysis when the substrate is covered with an initially incompressible film made auxetic by orthogonal oval-shaped voids, leading to an effective Poisson ratio ranging between -0.2 and 0 [9]. We see that the critical stretch ratio $\lambda_{c}\simeq 1.274$ matches well (within 5%) with the results in Fig. 4.
This paper’s analysis could play a critical role in manufacturing and testing the applicability of auxetic hydrogel organ patches [9] and skin grafts [49].
Appendix A: Validation of PBCs code – 2D incompressible neo-Hookean bilayer model
Figure A.1: (a) Schematic representation of an incompressible neo-Hookean stiff film/soft substrate bilayer system under uni-axial compression. On the left edge of the domain, displacement and shear traction are set at zero. On the bottom edge of the domain, roller support restricts vertical displacement and shear traction, and the top surface of the film is traction-free. (b) Linear buckling solutions with critical strains $\varepsilon_{c}$ when ${\mu_{f}}/{\mu_{s}}=30,1000$, in line with the predictions of Cao and Hutchinson [42].
Under uni-axial, plane-strain compression, the theoretical critical strain is $\varepsilon_{c}=\nicefrac{{1}}{{4}}\left({3\mu_{s}}/{\mu_{f}}\right)^{\nicefrac{{2}}{{3}}}$ [42].
In Fig. A.1, we applied PBCs on the left and right edges of the domain, roller support on the bottom edge of the substrate, perfect bonding between film and substrate layers, and traction-free condition on the top surface of the film in the two cases ${\mu_{f}}/{\mu_{s}}=30,1000$. We find $\varepsilon_{c}=0.053,0.0052$, respectively, matching well with the theoretical solutions.
Having validated our 1D PBCs script along the X-direction on a 2D model, we then extended the PBCs to two dimensions, along the X- and Z-directions, to perform a linear buckling analysis on 3D bilayer systems under uniaxial tension.
The height ratio of substrate to film layers was taken as ${h_{{s}}}/{h_{{f}}}=163$, with the width of layers being 30.457 units. We modelled both film and substrate with the incompressible neo-Hookean model.
We used a hybrid 8-node plane strain quadrilateral element with quadratic interpolation and reduced integration (CPE8RH) for both film and substrate layers. We also applied periodic boundary conditions on the left and right edges of the domain. The minimum size of the mesh element is lower than the height of the film. We took 12,800 elements, and the numerical results converged and were consistent with the theoretical solutions.
Appendix B: Coefficients in asymptotic expressions
$\begin{split}c_{4}&=\frac{\pazocal{C}\left(45+\nu_{s}c_{4a}+\nu_{f}c_{4b}+\nu_{f}^{2}c_{4c}\right)}{20\times 2^{1/3}\times 3^{2/3}(\nu_{f}-\nu_{s})^{2}(-1+\nu_{s})(-3+4\nu_{s})^{4/3}},\\ c_{5}&=\frac{\pazocal{C}^{2}(-1+\nu_{f})(-1+2\nu_{s})(-3-4\nu_{f}+4\nu_{s})}{2^{2/3}\times 3^{1/3}(\nu_{f}-\nu_{s})^{2}(-3+4\nu_{s})^{5/3}},\\ c_{6}&=\frac{4725-2\nu_{f}^{4}c_{6a}-\nu_{s}c_{6b}+\nu_{f}^{3}c_{6c}+\nu_{f}^{2}c_{6d}+2\nu_{f}c_{6e}}{12600(3-4\nu_{s})^{2}(-1+\nu_{s})^{2}(-\nu_{f}+\nu_{s})^{3}},\\ \pazocal{C}&=\left((1-\nu_{f})(\nu_{s}-1)\right)^{1/3},\end{split}$ (B.1a)
and
$\begin{split}c_{4a}&=-64+(13-4\nu_{s})\nu_{s},\\ c_{4b}&=-71+\nu_{s}(66+(51-26\nu_{s})\nu_{s}),\\ c_{4c}&=56-92\nu_{s}+26\nu_{s}^{2},\\ c_{6a}&=33727+2\nu_{s}(-73474+\nu_{s}(122601\\ &+\nu_{s}(-92584+26681\nu_{s}))),\\ c_{6b}&=4410+\nu_{s}(82574+\nu_{s}(-288226\\ &+\nu_{s}(399999-260716\nu_{s}+67034\nu_{s}^{2}))),\\ c_{6c}&=95848+2\nu_{s}(-138872+\nu_{s}(46253\\ &+\nu_{s}(235988+\nu_{s}(-297317+106724\nu_{s})))),\\ c_{6d}&=-3509+\nu_{s}(-202720+\nu_{s}(841376+\nu_{s}(-1218634\\ &+\nu_{s}(643241+2(22435-53362\nu_{s})\nu_{s})))),\\ c_{6e}&=-11970+\nu_{s}(78479+\nu_{s}(-137927+\nu_{s}(-12652\\ &+\nu_{s}(271583+\nu_{s}(-276527+89714\nu_{s}))))).\end{split}$ (B.1b)
Similarly, the coefficients of higher-order terms in the asymptotic expression for critical wavenumber (Eq. 4) are
$\begin{split}d_{3}&=\frac{4+2\nu_{s}-11\nu_{s}^{2}+\nu_{f}\left(11-32\nu_{s}+26\nu_{s}^{2}\right)}{15\left(3-7\nu_{s}+4\nu_{s}^{2}\right)},\\ d_{4}&=\frac{\left(\frac{2}{3}\right)^{2/3}\pazocal{C}(1+2\nu_{f})(-1+2\nu_{s})}{(-3+4\nu_{s})^{4/3}},\\ d_{5}&=\frac{\pazocal{C}^{2}\left(-6761+\nu_{s}d_{5a}+4\nu_{f}^{2}d_{5b}-2\nu_{f}d_{5c}\right)}{3150\times 2^{2/3}\times 3^{1/3}(-1+\nu_{f})(-1+\nu_{s})^{3}(-3+4\nu_{s})^{5/3}},\end{split}$ (B.2a)
and
$\begin{split}d_{5a}&=21724+\nu_{s}(-19146+\nu_{s}(-2636+7519\nu_{s})),\\ d_{5b}&=646+\nu_{s}(-4964+\nu_{s}(12906\\ &+\nu_{s}(-14204+5791\nu_{s}))),\\ d_{5c}&=-3821+\nu_{s}(7864+\nu_{s}(5844\\ &+\nu_{s}(-22796+13609\nu_{s}))).\end{split}$ (B.2b)
[…]
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 956401.
Supplementary data
There is no Supplementary material attached to this article.